OpenStudy (anonymous):

I have this question for homework but I don't understand this one bit. I did 9 out of 10 questions so this was the only one was stuck on... Could you please explain it to me. What is the mass, in grams, of a sample of 1.20 × 1022 atoms of mercury (Hg)? Show your work or explain the steps that you used to determine your answer.

OpenStudy (anonymous):

1.20 X 10.22

OpenStudy (anonymous):

@telijahmed can I have some help please

OpenStudy (anonymous):

@thomaster could I have some help please

OpenStudy (anonymous):

This is what I got but not sure if I did this right. 1.20 * 10^22 atoms Hg / 6.02 * 10^23 atoms / mole = 0.0199 moles Hg. atomic weight Hg = 200.6 g/mol. mass Hg = moles * at wt = 0.0199 mol * 200.6 g/mol = 3.99 g.

thomaster (thomaster):

\(\bf\Huge Welcome~to~\color{#00B4ff}{Open}\color{#7cc517}{Study}~!!!\) Let's have a look at the attached diagram.

OpenStudy (anonymous):

Wow thats a confusing diagram, I'm not sure if I understand.

thomaster (thomaster):

First divide by avogadro's number. \(\Large \dfrac{1.20 * 10^{22}}{6.022 * 10^{23}}=0.0199\ mol\)

OpenStudy (anonymous):

okay, thats what I got too.

thomaster (thomaster):

Then multiply with molar mass (200.6 g/mol) 0.0199 * 200.6 = 3.997 gram

OpenStudy (anonymous):

so would the answer be 3.997 grams

thomaster (thomaster):

yes :)

OpenStudy (anonymous):

So was my original answer above correct.

thomaster (thomaster):

Yes it was correct. I only calculated it with more decimals. By the way the diagram is not that confusing. You start at nr of particles. Follow the arrow to mole. It says divide by avogadro's number, then you have the moles. Last step is from moles to mass, so multiply with molar mass :)

OpenStudy (anonymous):

okay thank you for your help, much appreicated!

thomaster (thomaster):

You're welcome